# Unifying 2D and 3D Vision-Language Understanding

### v2

@@ 1 Introduction
Today’s real-world embodied systems rely on depth sensors and egocentric, calibrated camera setups for navigation and interaction with their surroundings (Ahn et al., 2022; Chiang et al., 2024). However, despite having access to rich 3D information, these systems predominantly use 2D vision-language models to interpret their sensory video input, rather than leveraging 3D models that incorporate depth and egomotion. At first glance, this reliance on 2D models appears counterintuitive, as prior research has consistently shown that 3D models outperform their 2D counterparts when trained on comparable amounts of data (Siddiqui et al., 2023; Kundu et al., 2020a; Fang et al., 2021; Rukhovich et al., 2022). The key limitation, however, is dataset availability: while 2D datasets are vast and well-curated, 3D datasets remain scarce and expensive to annotate (Dai et al., 2017; Yeshwanth et al., 2023). As a result, there are currently no high-performing, pre-trained 3D encoders capable of processing 3D inputs at the same level as CLIP (Radford et al., 2021) does for 2D images. This data imbalance has led to a significant performance gap, ultimately slowing the widespread adoption of 3D models in embodied vision-language systems. Given these challenges, is scaling 3D training data the only viable path to bridging this gap, or are there alternative strategies for making 3D models more effective?
@@ 2 Related Work
For additional related work, see Appendix (Section-A.14).
@@ 3.1 Supervision Objective
Implementation details: UniVLG consists of 108M trainable parameters along with a frozen 220M parameter text-encoder (Koukounas et al., 2024) and a 304M parameter image-encoder (Oquab et al., 2024). At test time, we feed all images to our model with a 90-frame scene taking $\sim$1050ms and $\sim$15GB of VRAM on an A100 GPU. We refer to Section A.2 for further details.
@@ 4 Experiments
UniVLG-3D-only (Ours) 71.6 63.8 49.4 78.9 52.5 43.3 34.2 65.8 60.7 53.2 42.6
UniVLG (Ours) 73.0 64.8 51.8 - 56.3 48.0 37.6 - 63.5 56.4 46.0
@@ 4.1 Evaluation on 3D Referential Grounding
Table 2: Out-of-Domain 3D Referential Grounding Acc@25 in Det. From left-to-right, ScanNet++, HM3D, ARKitScenes, ScanNet (GT), ScanNet (SAMPro3D). See Section A.3 for details.
@@ 4.2 Evaluation on Out-of-Domain 3D Referential Grounding
We evaluate our model and baselines on LX3Dfootnotetext: LX3D is contributed by a concurrent anonymous submission and is not part of our contributions., an out-of-domain 3D language grounding dataset that spans ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), ARKitScenes (Baruch et al., 2021), and HM3D (Yadav et al., 2023) (details in Section A.3). LX3D allows us to assess the robustness of our model on new scenes, camera capture systems, and language instructions. We show the results of our model, both a 3D-only variant and our full model w/2D data + lifting in Table 2. We find that our model outperforms prior methods on these out-of-domain datasets, achieving strong performance across the board.
@@ 4.3 Evaluation on 3D Question Answering
Baselines: We compare against the LLM based methods of 3D-LLM (Hong et al., 2023) and NaviLLM (Zheng et al., 2024) which use BLIP2-flanT5 (Li et al., 2023) and Vicuna-7B (Peng et al., 2023a) as their answer generation heads. We also compare with 3D-VisTA (Zhu et al., 2023b) and PQ3D (Zhu et al., 2024b) which use small decoder heads like T5-small (Raffel et al., 2020), similar to our approach. We show results in Table 3 on the validation sets of these benchmarks. UniVLG outperforms all prior baselines on both benchmarks. We found that using sensor point clouds vs mesh point clouds does not result in a significant difference in performance in these benchmarks, likely because the models are evaluated on text generation instead of localization of objects as in 3D referential grounding and segmentation benchmarks.
Additionally, we show results on 3D instance segmentation in Section A.1.
@@ 4.5 Ablations
1. Lifting 2D datasets to 3D improves 3D performance. In Table 6, we compare three variants of our model: one trained only on 3D data, one trained with 3D data and 2D images without lifting them to 3D (where the 3D layers are skipped for 2D inputs, following (Jain et al., 2024)), and our proposed approach of lifting 2D images to 3D pointmaps. We observe that incorporating 2D data improves performance in both scenarios, but our approach of lifting 2D images to 3D achieves the best results. In Section A.11, we show in a more controlled setting that training without using 3D pointmaps—by skipping the 3D layers—results in significant overfitting to individual 2D-3D domains.
@@ 5 Conclusion
6 Acknowledgements
@@ A.1 Evaluation on 3D Instance Segmentation
We test UniVLG on 3D segmentation benchmarks of ScanNet200 (Rozenberszki et al., 2022) and Matterport3D (Chang et al., 2017) for instance segmentation tasks. These benchmarks have a fixed vocabulary of objects (200 classes in ScanNet200 and 160 classes in Matterport3D). SOTA models like ODIN (Jain et al., 2024) and Mask3D (Schult et al., 2023) train and evaluate in this fixed vocabulary setup by predicting a distribution over the fixed set of classes and supervising with softmax losses. PQ3D (Zhu et al., 2024b) evaluates in a language-prompted setup where they supply object names, one object at a time, and gather predictions for all objects in the vocabulary. They compare with a closed-vocabulary version of their model, and find that their language-prompted version is about 7% worse than their closed vocabulary version due to ambiguities in class names confusing CLIP (eg. “chair" and “armchair"; “table" and “desk" are different categories in ScanNet200). We follow PQ3D and evaluate our model in the language-prompted setup. The input to the model is a concatenation of all object classes of the benchmark as a long sentence (eg: “chair. table. sofa. bed. …."). While PQ3D cannot predict multiple object classes simultaneously, and hence have to supply one object at a time, our model can simulatenously decode masks for all objects mentioned in the sentence. The results are shown in Table-8 on the official validation splits of these benchmarks. We observe that UniVLG outperforms PQ3D in the language-prompted evaluation setup on ScanNet200.
A.2 Additional Implementation details
UniVLG consists of 108M trainable parameters along with a frozen 220M parameter text-encoder (Koukounas et al., 2024) and a 304M parameter image-encoder (Oquab et al., 2024). For ablations in Table 7 and 5, we use a 88M parameter Swin (Liu et al., 2021) image-encoder. We train in data-parallel across 32 A100 80G GPUs with an effective batch size of 64. We use ScanEnts3D (Abdelreheem et al., 2023) version of ScanRefer (Chen et al., 2020a) and Referit3D (Achlioptas et al., 2020) which provides object annotations for all noun words in the language sentence. During training, we process either a sequence of $N$ posed RGB-D images, or a single RGB image. During training, the model processes either a sequence of $N$ posed RGB-D images or a single RGB image. For 2D images, we apply a 2D-to-3D lifting strategy with a 50% probability. When lifted, the images pass through all 2D-3D layers; otherwise, they remain in 2D space, skipping the 3D attention layers. At test time, we retain 2D images in their original space to prevent noise from predicted 3D pointmaps from impacting 2D performance. For 3D scenes, we compute CLIP embeddings for all images and captions and use this to select 5 relevant frames, with an additional 10 frames coming from Furthest-Point-Sampling (FPS) in the CLIP embedding space, for a total of 15 frames. At test time, we feed all images in a scene to our model. For validation results, we perform span prediction to identify the primary subject from a given utterance. We found that prompting an LLM (Dubey et al., 2024) with examples specific to a given dataset to result in better performance compared to traditional NLP libraries. We use Jina-CLIP (Koukounas et al., 2024) as the text-encoder, as it supports arbitrary input-length. We jointly train our model on all datasets, with text generation loss only active in question answering datasets. Our method provides for fast inference, with a 90-frame scene taking $\sim$1050ms and $\sim$15GB of VRAM on an A100 GPU.
A.3 More details on Out-of-domain LX3D datasets
Note: This collection of these out-of-domain datasets is not our contribution, but comes from a concurrent anonymous submission. We include the following details for the information of the reader.
The data is collected using an off-the-shelf annotation interface which allows for selecting rectangular regions on 2D videos and associating them with text. 3D instance masks are projected into 2D and shown alongside the RGB camera view. For ScanNet-GT (Dai et al., 2017), HM3D (Yadav et al., 2023) and ScanNet++ (Yeshwanth et al., 2023), ground truth instance masks are used that are provided by the dataset. For ARKitScenes (Baruch et al., 2021), and ScanNet-SAM, instance masks produced by SAMPro3D (Xu et al., 2023) are used. Annotators have the ability to conglomerate multiple instance masks into a single object – this proved necessary for SAMPro3D segmentations, which tended to break large objects into many subparts. This is in line with ScanNet’s instance segmentation procedure in which annotators stitch together an automatically-generated oversegmentation of a 3D scene.
In addition to grounding the target object, annotators were asked to provide object grounding for all other nouns in the phrase they provided. For SAMPro3D masks, annotators could select any object they wished as a target. In later iterations of the task, for ScanNet and ScanNet++, a mask was highlighted in white with at least 1 distractor to serve as the target object for description. This is in line with prior literature which observed that the 3D referential grounding task is most challenging in the high-distractor case (Chen et al., 2020a).
While it was found that annotations on ScanNet and ScanNet++ were relatively reliable from a first pass, considerable mask/bounding box noise was observed when collecting annotations on SAMPro3D instance masks. To address this, a separate validation task was created. In this task, annotators saw a video similar to the video provided in the generation task, but only the selected masks for a particular language sample were highlighted. The referential expression description was also shown. This improved the quality of SAMPro3D samples substantially, but it also drastically slowed the collection rate.
A.3.1 Annotation setup
The annotation setup largely mirrors ScanRefer (Chen et al., 2020a) – annotators simply describe objects in a one-step fashion rather than the two-player game format of NR3D (Achlioptas et al., 2020). A 2D video interface was used for the task, projecting 3D instance masks to 2D. Annotators can write full referring expressions and associate particular instance masks with particular tokens in the expressions. They may also agglomerate two or more instances into one.
When ground-truth instance masks are available (as in ScanNet and ScanNet++), these are used as the basis for annotation. When they are not available, instance masks are generated using SAMPro3D (Xu et al., 2023).
Decomposed by scene dataset, LX3D contains:
  1. 1.
    ScanNet: 4,470 language annotations covering 130 venues and 1038 objects for validation.
  2. 2.
    ScanNet++: 3,774 language annotations covering 50 venues and 1,303 objects for validation.
  3. 3.
    ARKitScenes: 1,416 language annotations covering 714 venues and 2,322 objects for validation.
A.4 Effect of Fine-tuning 2D backbones in UniVLG
@@
We study the effect of fine-tuning the 2D backbones on in-domain and out-of-domain performance. We train two versions of UniVLG with swin backbones, one with fine-tuning and the other without fine-tuning. For training, we use SR3D and NR3D, and evaluate on the validation sets of SR3D, NR3D (in-domain) and ScanRefer (out-of-domain). The results of the experiments are shown in Table-9. We find that both models work similarly well, both in-domain and out-of-domain.
@@
A.5 Performance with different backbones
We demonstrate that the performance can scale with the strength of the backbone. Specifically we use a DINOv2 (Oquab et al., 2024) backbone consisting of 1.1B parameters, scaling over 5x compared to the Swin backbone. To achieve high-performance during training, we freeze the backbone, although we note that it is possible that additional performance could be obtained with efficient fine-tuning techniques such as LoRA (Hu et al., 2021). In Table-10, we find that adding this backbone boosts performance on all 3 language grounding datasets, with substantial margins of $4.5\%$, $1.9\%$, and $3.4\%$ 0.25 on SR3D, NR3D, and ScanRefer respectively.
@@ A.5 Performance with different backbones
A.6 Additional Metrics on ScanQA Dataset
We report additional standard metrics used by ScanQA benchmark in Table-11.
A.7 Visualizations of UniVLG on Referential Grounding Datasets
@@ A.8 Visualization of common failure modes of UniVLG
A.9 Performance analysis with pose and depth noise
@@ A.10 Miscellaneous Details
A.11 2D-3D Generalization Test
Table 12: 2D-3D Generalization Test
Model 3D Sup. Classes 2D Sup. Classes
UniVLG 72.6 53.8
UniVLG w/o 2D-to-3D lifting 71.4 0.0
UniVLG (Upper-Bound) 69.7 84.2
We study the generalization between 2D and 3D domains using the 3D Instance Segmentation benchmark in the AI2-THOR simulator (Kolve et al., 2017). AI2-THOR consists of 120 object classes and provides posed RGB-D images of its 3D environments. We split these classes into two disjoint subsets: (a) 3D Supervised Classes (30) and (b) 2D Supervised Classes (90).
We train UniVLG in a 2D-3D setting where, for multi-view RGB-D inputs, the model is tasked with detecting objects from the 3D Supervised Classes subset. For single RGB image inputs, it is tasked with detecting objects from the 2D Supervised Classes subset. At test time, the model is evaluated in a multi-view posed RGB-D setup on all 120 object classes to assess its generalization across the two subsets. Results are presented in Table 12.
We find that when UniVLG does not lift 2D images to 3D—simply skipping 3D attention layers in 2D batches, similar to ODIN (Jain et al., 2024) (row 2 in Table 12)—performance on 2D Supervised Classes drops to nearly zero. This suggests that without depth-based transformation, the model fails to learn from 2D supervision. However, when UniVLG incorporates depth information, its performance on 2D Supervised Classes becomes non-trivial, indicating that it successfully learns from 2D supervision and applies this knowledge in the 3D domain at test time.
In row 3, we report results for UniVLG trained in 3D on all 120 classes, representing an upper bound on its performance given full 3D supervision. We observe that UniVLG in row 1 remains significantly below this upper bound, highlighting the need for further improvements in 2D-to-3D generalization.
A.12 Discussion: Decoding Masks vs. Boxes
@@
A.13 Evaluation with Mask based Intersection Over Union

### v3

@@ 1 Introduction
Today’s real-world embodied systems rely on depth sensors and egocentric, calibrated camera setups for navigation and interaction with their surroundings (Ahn et al., 2022; Chiang et al., 2024). However, despite having access to rich 3D information, these systems predominantly use 2D vision-language models to interpret their sensory video input, rather than leveraging 3D models that incorporate depth and egomotion. At first glance, this reliance on 2D models appears counterintuitive, as prior research has consistently shown that 3D models outperform their 2D counterparts when trained on comparable amounts of data (Siddiqui et al., 2023; Kundu et al., 2020a; Fang et al., 2021; Rukhovich et al., 2022). The key limitation, however, is dataset availability: while 2D datasets are vast and well-curated, 3D datasets remain scarce and expensive to annotate (Dai et al., 2017; Yeshwanth et al., 2023). As a result, there are currently no high-performing, pre-trained 3D encoders capable of processing 3D inputs at the same level as CLIP (Radford et al., 2021) does for 2D images. This data imbalance has led to a significant performance gap, ultimately slowing the widespread adoption of 3D models in embodied systems. Given these challenges, is scaling 3D training data the only viable path to bridging this gap, or are there alternative strategies for making 3D models more effective?
@@ 2 Related Work
For additional related work, see Appendix (Section-A.10).
@@ 3.1 Supervision Objective
Implementation details: UniVLG consists of 108M trainable parameters along with a frozen 220M parameter text-encoder (Koukounas et al., 2024) and a 304M parameter image-encoder (Oquab et al., 2024). For ablations in Table 7 and 5, we use a 88M parameter Swin (Liu et al., 2021) image-encoder. We train in data-parallel across 32 A100 80G GPUs with an effective batch size of 64. We use ScanEnts3D (Abdelreheem et al., 2023) version of ScanRefer (Chen et al., 2020a) and Referit3D (Achlioptas et al., 2020) which provides object annotations for all noun words in the language sentence. During training, we process either a sequence of $N$ posed RGB-D images, or a single RGB image. During training, the model processes either a sequence of $N$ posed RGB-D images or a single RGB image. For 2D images, we apply a 2D-to-3D lifting strategy with a 50% probability. When lifted, the images pass through all 2D-3D layers; otherwise, they remain in 2D space, skipping the 3D attention layers. At test time, we retain 2D images in their original space to prevent noise from predicted 3D pointmaps from impacting 2D performance. For 3D scenes, we compute CLIP embeddings for all images and captions and use this to select 5 relevant frames, with an additional 10 frames coming from Furthest-Point-Sampling (FPS) in the CLIP embedding space, for a total of 15 frames. At test time, we feed all images in a scene to our model. For validation results, we perform span prediction to identify the primary subject from a given utterance. We found that prompting an LLM (Dubey et al., 2024) with examples specific to a given dataset to result in better performance compared to traditional NLP libraries. We use Jina-CLIP (Koukounas et al., 2024) as the text-encoder, as it supports arbitrary input-length. We jointly train our model on all datasets, with text generation loss only active in question answering datasets. Our method provides for fast inference, with a 90-frame scene taking $\sim$1050ms and $\sim$15GB of VRAM on an A100 GPU.
@@ 4 Experiments
UniVLG-3D-only (Ours) 71.6 63.8 49.4 81.7 54.7 44.9 35.7 65.2 60.7 53.2 42.6
UniVLG (Ours) 73.0 64.8 51.8 - 58.3 49.8 39.1 - 63.5 56.4 46.0
@@ 4.1 Evaluation on 3D Referential Grounding
Table 2: Out-of-Domain 3D Referential Grounding Acc@25 in Det. From left-to-right, ScanNet++, HM3D, ARKitScenes, ScanNet (GT), ScanNet (SAMPro3D). See LABEL:appendix:ood_data for details.
@@ 4.2 Evaluation on Out-of-Domain 3D Referential Grounding
We evaluate our model and baselines on L3DD (Arnaud et al., 2025), an out-of-domain 3D language grounding dataset that spans ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), ARKitScenes (Baruch et al., 2021), and HM3D (Yadav et al., 2023). L3DD allows us to assess the robustness of our model on new scenes, camera capture systems, and language instructions. We show the results of our model, both a 3D-only variant and our full model w/2D data + lifting in Table 2. We find that our model outperforms prior methods on these out-of-domain datasets, achieving strong performance across the board.
@@ 4.3 Evaluation on 3D Question Answering
Baselines: We compare against the LLM based methods of 3D-LLM (Hong et al., 2023) and NaviLLM (Zheng et al., 2024) which use BLIP2-flanT5 (Li et al., 2023) and Vicuna-7B (Peng et al., 2023a) as their answer generation heads. We also compare with 3D-VisTA (Zhu et al., 2023b) and PQ3D (Zhu et al., 2024b) which use small decoder heads like T5-small (Raffel et al., 2020), similar to our approach. We show results in Table 3 on the validation sets of these benchmarks. UniVLG outperforms all prior baselines on both benchmarks. We found that using sensor point clouds vs mesh point clouds does not result in a significant difference in performance in these benchmarks, likely because the models are evaluated on text generation instead of localization of objects as in 3D referential grounding and segmentation benchmarks. Additionally, we show results on 3D instance segmentation in Section A.1.
@@ 4.5 Ablations
1. Lifting 2D datasets to 3D improves 3D performance. In Table 6, we compare three variants of our model: one trained only on 3D data, one trained with 3D data and 2D images without lifting them to 3D (where the 3D layers are skipped for 2D inputs, following (Jain et al., 2024)), and our proposed approach of lifting 2D images to 3D pointmaps. We observe that incorporating 2D data improves performance in both scenarios, but our approach of lifting 2D images to 3D achieves the best results. In Section 4.6, we show in a more controlled setting that training without using 3D pointmaps—by skipping the 3D layers—results in significant overfitting to individual 2D-3D domains.
@@
4.6 2D-3D Generalization Test
Table 8: 2D-3D Generalization Test
Model 3D Sup. Classes 2D Sup. Classes
UniVLG 72.6 53.8
UniVLG w/o 2D-to-3D lifting 71.4 0.0
UniVLG (Upper-Bound) 69.7 84.2
We study the generalization between 2D and 3D domains using the 3D Instance Segmentation benchmark in the AI2-THOR simulator (Kolve et al., 2017). AI2-THOR consists of 120 object classes and provides posed RGB-D images of its 3D environments. We split these classes into two disjoint subsets: (a) 3D Supervised Classes (30) and (b) 2D Supervised Classes (90). We train UniVLG in a 2D-3D setting where, for multi-view RGB-D inputs, the model is tasked with detecting objects from the 3D Supervised Classes subset. For single RGB image inputs, it is tasked with detecting objects from the 2D Supervised Classes subset. At test time, the model is evaluated in a multi-view posed RGB-D setup on all 120 object classes to assess its generalization across the two subsets. Results are presented in Table 8.
We find that when UniVLG does not lift 2D images to 3D—simply skipping 3D attention layers in 2D batches, similar to ODIN (Jain et al., 2024) (row 2 in Table 8)—performance on 2D Supervised Classes drops to nearly zero. This suggests that without depth-based transformation, the model fails to learn from 2D supervision. However, when UniVLG incorporates depth information, its performance on 2D Supervised Classes becomes non-trivial, indicating that it successfully learns from 2D supervision and applies this knowledge in the 3D domain at test time. In row 3, we report results for UniVLG trained in 3D on all 120 classes, representing an upper bound on its performance given full 3D supervision. We observe that UniVLG in row 1 remains significantly below the upper bound, highlighting the need for further improvements in 2D-to-3D generalization.
4.7 Common failure modes of UniVLG
We identify three systematic failure modes in our model, illustrated in Figure-5 (see Appendix).
Inclusion of distant outlier points in the predicted masks: In the first image of Figure-5, while UniVLG accurately predicts the object, it also includes some distant points in the mask. This leads to a larger bounding box during the mask-to-bounding box conversion in post-processing, negatively affecting accuracy metrics. Although our proposed box loss mitigates this issue, it doesn’t fully resolve it.
Multiple instances of the same object being segmented together: As shown in the middle image of Figure-5, UniVLG predicts both beds as a single output. Incorporating attention to language and queries and box loss helps reduce such errors, though they still persist.
Grounding failures as seen in the third image of Figure-5.
@@ 5 Conclusion
While our current approach already achieves strong performance using significantly less 2D data than state-of-the-art 2D VLMs, scaling to larger 2D datasets is a promising next step to further improve 3D reasoning. Additionally, although UniVLG is designed for static 3D scenes, extending it to dynamic 3D environments remains an important and exciting direction for future work. Finally, integrating UniVLG with large-scale vision-language models holds strong potential for unifying 2D and 3D understanding at scale.
Acknowledgements
@@ A.1 Evaluation on 3D Instance Segmentation
We test UniVLG on 3D segmentation benchmarks of ScanNet200 (Rozenberszki et al., 2022) and Matterport3D (Chang et al., 2017) for instance segmentation tasks. These benchmarks have a fixed vocabulary of objects (200 classes in ScanNet200 and 160 classes in Matterport3D). SOTA models like ODIN (Jain et al., 2024) and Mask3D (Schult et al., 2023) train and evaluate in this fixed vocabulary setup by predicting a distribution over the fixed set of classes and supervising with softmax losses. PQ3D (Zhu et al., 2024b) evaluates in a language-prompted setup where they supply object names, one object at a time, and gather predictions for all objects in the vocabulary. They compare with a closed-vocabulary version of their model, and find that their language-prompted version is about 7% worse than their closed vocabulary version due to ambiguities in class names confusing CLIP (eg. “chair" and “armchair"; “table" and “desk" are different categories in ScanNet200). We follow PQ3D and evaluate our model in the language-prompted setup. The input to the model is a concatenation of all object classes of the benchmark as a long sentence (eg: “chair. table. sofa. bed. …."). While PQ3D cannot predict multiple object classes simultaneously, and hence have to supply one object at a time, our model can simulatenously decode masks for all objects mentioned in the sentence. The results are shown in Table-9 on the official validation splits of these benchmarks. We observe that UniVLG outperforms PQ3D in the language-prompted evaluation setup on ScanNet200.
A.2 Effect of Fine-tuning 2D backbones in UniVLG
@@
We study the effect of fine-tuning the 2D backbones on in-domain and out-of-domain performance. We train two versions of UniVLG with swin backbones, one with fine-tuning and the other without fine-tuning. For training, we use SR3D and NR3D, and evaluate on the validation sets of SR3D, NR3D (in-domain) and ScanRefer (out-of-domain). The results of the experiments are shown in Table-10. We find that both models work similarly well, both in-domain and out-of-domain.
@@
A.3 Performance with different backbones
We demonstrate that the performance can scale with the strength of the backbone. Specifically we use a DINOv2 (Oquab et al., 2024) backbone consisting of 1.1B parameters, scaling over 5x compared to the Swin backbone. To achieve high-performance during training, we freeze the backbone, although we note that it is possible that additional performance could be obtained with efficient fine-tuning techniques such as LoRA (Hu et al., 2021). In Table-11, we find that adding this backbone boosts performance on all 3 language grounding datasets, with substantial margins of $4.5\%$, $1.9\%$, and $3.4\%$ 0.25 on SR3D, NR3D, and ScanRefer respectively.
@@ A.3 Performance with different backbones
A.4 Additional Metrics on ScanQA Dataset
We report additional standard metrics used by ScanQA benchmark in Table-12.
A.5 Visualizations of UniVLG on Referential Grounding Datasets
@@
A.6 Performance analysis with pose and depth noise
@@ A.7 Miscellaneous Details
A.8 Discussion: Decoding Masks vs. Boxes
@@
A.9 Evaluation with Mask based Intersection Over Union